What is the concentration of stock HNO3 which has density 1.50gcm-3 and has purity of 70%


1
Expert's answer
2021-09-16T02:01:17-0400

Molar mass MM(HNO3) = 63 g/mol

Density ρ = 1.42 g/ml

Purity = 70 %

Molarity

"M = \frac{m}{MM \times V} \\

ρ = \frac{m}{V} \\

M = \frac{ρ \times 1000}{MM} \times Purity \\

M = \frac{1.50 \times 1000}{63} \times \frac{70}{100} = 16.66 \;moles/L"

Answer: 16.66 M